OpenStudy (anonymous):

Please help with some Synthetic Division!!! 6x^3-25x^2+2x+8/2x+1

OpenStudy (anonymous):

anything?

OpenStudy (anonymous):

hold on i might be able to help

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

The rational root theorem for polynomials with integer coefficients states that the numerator of any rational root must divide the constant term, and the denominator must divide the coefficient of the highest power. So guessing various 2x2x2's over 2x3's we get lucky at x=4. The next step is to divide (x-4) out of the original polynomial to reduce the remaining part to a quadratic: 6x^3 – 25x^2 + 2x + 8 = ( x - 4 ) ( 6x^2 - x - 2 ) Then use whatever favorite method you have for the quadratic roots. I always just start with the quadratic formula since it handles all cases in a snap: x = [ 1 +/- sqrt( 1 + 48 ) ] / 12 = [ 1 +/- 7 ] / 12 = { 2/3, -1/2 } The complete factorization is therefore ( x - 4 ) ( 3x - 2 ) ( 2x + 1 )

OpenStudy (anonymous):

dont hate me if i get this wrong

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i dont understand why we divide by 4??

OpenStudy (anonymous):

hold up i answerd it wrong bcz i didnt totally understand it but i think i get it now give me a sec

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

damn im so sorry im lost i cant help

OpenStudy (anonymous):

fuk u

OpenStudy (anonymous):

wtf u shouldnt be on this site then

OpenStudy (anonymous):

stupid retrice

OpenStudy (anonymous):

ur a real penis hole